OpenStudy (megannicole51):

A metal plate of constant density has a shape bounded by the curve y=sqrt(x), the x-axis, and the line x=1. Please help trying to study for exam tomorrow!

OpenStudy (ranga):

The bounded area looks like this. The enclosed area is the area of interest. 

OpenStudy (megannicole51):

okay:D got that so far!

OpenStudy (kmeis002):

Are you trying to find the weight?

OpenStudy (ranga):

There is no question in the problem. Just a statement!

OpenStudy (megannicole51):

Oh im sorry...im just trying to find the mass of the plate and the center of the mass

OpenStudy (ranga):

You need to integrate the function y = sqrt(x) between the limits x = 0 and x = 1 to get the area under the curve. Assume the plate has a constant thickness t and density d then Volume = Area x t Mass = Density x Volume = d x Area x t Substitute Area from the integration.

OpenStudy (megannicole51):

integration of sqrt(x)=(2/3)

OpenStudy (megannicole51):

hey @ranga i just looked at the answer sheet and apparently they messed up and didnt add the density....so the constant density is 5

OpenStudy (megannicole51):

so the mass im good with:) its the center mass that i am confused with....i know that you have to either have x, y, or z when you are given the constant density, however, i only know how to do x

OpenStudy (ranga):

Oh, the density is given as mass per surface area and not as mass per volume. So you can ignore the thickness t.

OpenStudy (megannicole51):

yeah i realized that lol:)

OpenStudy (ranga):

There is a standard formula for finding the x and y coordinates of the center of mass which is what they are using.

OpenStudy (megannicole51):

do u know what they are?

OpenStudy (megannicole51):

and x im a little confused on how they are getting their answer...like i know how to set it up but solving it is where i get mixed up

OpenStudy (ranga):

\[X = \frac{ 1 }{ A } \int\limits_{a}^{b}xf(x)dx\] X is x coordinate of center of mass and A is the area of the enclosed region found earlier.

OpenStudy (ranga):

\[Y = \frac{ 1 }{ A}\int\limits_{a}^{b}\frac{ 1 }{ 2}(f(x))^{2}dx\]

OpenStudy (megannicole51):

so can you show out the steps to x and y for me so i can see it being done please?

OpenStudy (ranga):

\[A = \int\limits_{a}^{b}f(x)dx\]

OpenStudy (megannicole51):

 so this is where i get stuck...